ganeshie8 (ganeshie8):

Find the remainder when \( 4^{2015} + 5^{2015}\) is divided by \(9\)

ganeshie8 (ganeshie8):

In other words, work/reduce : \[( 4^{2015} + 5^{2015}) \mod 9 \]

OpenStudy (anonymous):

using binomial 9=4+5. \[4^{2014} +5^{2014}= (4+5)[4^{2013} +5 *4^{2012}+.....+5^{2013}]\]

ganeshie8 (ganeshie8):

Clever ! if i understand correctly it seems you have used the formula \(x^n - y^n\) and factored 4+5 which shows the number is a multiple of 9. Interesting xD

OpenStudy (anonymous):

yea

hartnn (hartnn):

\(x^n+y^n\) and there will be alternate negatives right?? it won't affect the final result though :)

hartnn (hartnn):

any other methods ?

ganeshie8 (ganeshie8):

Oh right! thats a good catch, the terms in sum should alternate signs..

hartnn (hartnn):

\(x^n~ mod ~q\) there's any theorem ? when q is prime ?

ganeshie8 (ganeshie8):

im looking forward to see many more clever methods xD i didn't think of the catch.me's trick.. it was pretty neat :)

ganeshie8 (ganeshie8):

yes but 9 is not prime here so we need to try something else i think

hartnn (hartnn):

\(x^n~ mod ~ab = x^n~ mod~a \times x^n ~mod~b\) is something like that true ?

OpenStudy (mathmath333):

the last digit of \(5^n\) is always \(5\) so u need to work only \(4^{2015}\)

OpenStudy (mathmath333):

their is clever method lol

ganeshie8 (ganeshie8):

lets see.. consider \(2^3\pmod{9}\) \(2^3 \mod{3}\) gives 2 so \(2^3 \mod{3} \times 2^3 \mod{3} = 2 \times 2 = 4\) however we know that \(2^3\pmod{9}\) equals 8 so this relation will not hold in general @hartnn

hartnn (hartnn):

math, if 5^n last digit is 5, then we can ignore it in calculation mod 9 ? i didn't get that point

OpenStudy (mathmath333):

\(\large\tt \begin{align} \color{black}{ \dfrac{4^{2015} + 5^{2015}}{9}\\~\\ \implies \dfrac{(9-5)^{2015} + 5^{2015}}{9}\\~\\ \implies \dfrac{(-5)^{2015} + 5^{2015}}{9}\\~\\ \Large \equiv 0}\end{align}\)

ganeshie8 (ganeshie8):

Brilliant !!!!!!

OpenStudy (mathmath333):

oh i see yes the last digit only works on mod 10 sry

ganeshie8 (ganeshie8):

that really cute :D

hartnn (hartnn):

if shorter methods are done, post some longer traditional methods too :P

ganeshie8 (ganeshie8):

we can generalize this i think \(a^{2k+1} + b^{2k+1}\) is always divisible by \(a+b\)

OpenStudy (mathmath333):

@hartnn tell me remainder \(\dfrac{6^{159}}{(10)}\)

OpenStudy (anonymous):

i think you need the final equation no proof!

ganeshie8 (ganeshie8):

"traditional" reminds me of binomial theorem, it seems bit lengthy...

hartnn (hartnn):

6 is correct? math

OpenStudy (mathmath333):

yes how u got that

hartnn (hartnn):

last digit 6^1 =6 6^2= 6 6^3 =6 so on,

ganeshie8 (ganeshie8):

thats a fun problem haha! guess 6 is the only number that stays 6 after taking any number of powers

ganeshie8 (ganeshie8):

1 is boring : 1^1 = 1 1^2 = 1 1^3 = 1 ...

OpenStudy (mathmath333):

yes thast what i was saying if its mod 10 ,then the last digit trick is very powerful

hartnn (hartnn):

2 has a cycle of (2,4,8,6)

OpenStudy (mathmath333):

5^1 5^2..........has unit digit 5

ganeshie8 (ganeshie8):

Oh yes ! so 1, 5 and 6

OpenStudy (mathmath333):

9 has 9,1,9,1,91,9,1,9,1,9,1,9,1,9,1,9......

hartnn (hartnn):

1 -- (1) 2 -- (2,4,8,6) 3 -- (3, 9,7, 1) 4 -- (4, 6) 5 -- (5) 6 -- (6) 7 --(7,9,3,1) 8 -- (8,4,2,6) 9 -- (9,1) all correct ?

hartnn (hartnn):

i observe, whenever 1 comes in the cycle, thats the end...obviously :P

OpenStudy (mathmath333):

u missed 0 , lol

hartnn (hartnn):

0 -- (0) :P

hartnn (hartnn):

whenever 6 comes in the cycle, that is also the end... 6*2 =1 `2` 6*4 =2 `4` 6*6 = 3 `6` 6*8 =4 `8`

OpenStudy (anonymous):

Messy one \[4 \times4^{2014} +5\times 4^{2014}-5\times 4^{2014} +5\times5^{2014}+4\times5^{2014}-4\times 5^{2014}\] \[9\times4^{2014}-9\times4^{2013}+9\times5^{2014}-9\times5^{2013}.\] thus remainder is Zero

OpenStudy (mathmath333):

this is too long

OpenStudy (anonymous):

why long you just plus and minus what makes it a multiple of 9

OpenStudy (mathmath333):

i mean the length is long to write such things in exam

ganeshie8 (ganeshie8):

thats kinda cute :D i see what you're doing but the second line is not matching with the first line

OpenStudy (anonymous):

revise it again

ganeshie8 (ganeshie8):

\[\color{Red}{4 \times4^{2014} +5\times 4^{2014}}-5\times 4^{2014} +\color{blue }{5\times5^{2014}+4\times5^{2014}}-4\times 5^{2014}\]

ganeshie8 (ganeshie8):

\[\color{Red}{9 \times4^{2014} }-5\times 4^{2014} +\color{blue }{9\times5^{2014}}-4\times 5^{2014}\]

ganeshie8 (ganeshie8):

how do you get a 9 out of remaining two terms ?

OpenStudy (anonymous):

yea I made 20 =9 sorry :(

OpenStudy (anonymous):

can i use phi :P phi(9)=phi(3^2)=3^2-3=6 5^6=1 mod 9 5^2010=1 mod 9 5^5= 2 mod 9 5^2015=2 mod 9---1 4^6=1 mod 9 4^2010=1 mod 9 4^5=7 mod 9 4^2015=7 mod 9----2 from 1 +2 4^2015+5^2015=0 mod 9

ganeshie8 (ganeshie8):

Neat :)

ganeshie8 (ganeshie8):

but the winner is this method for keeping it short and simple :D http://gyazo.com/566f88e1d0c62946d936d758dd62aae2

OpenStudy (anonymous):

:P

OpenStudy (anonymous):

uhh... sorry for nitpicking, guys, but uhmm Catch.me's method is awesome, but his example was kind of faulty, even WITH the alternating signs. \[4^{2014} +5^{2014}= (4+5)[4^{2013} -5 *4^{2012}+.....-5^{2013}]\] is not true. This factoring method for \(\large x^n + y^n \) Only works if n is odd. For instance, say n is 4. \[\Large 4^4 + 5^4 = 256+625 = 881\] 881 is not even divisible by 9 which is 4+5

ganeshie8 (ganeshie8):

thats right! we can factor like that only when the exponent is odd : \[4^{2015} +5^{2015}= 4^{2015} -(-5)^{2015} = (4+5)[stuff]\]

OpenStudy (anonymous):

Good thing 2015 is odd, yeah? haha

OpenStudy (anonymous):

speaking of 2015, happy new year, Ganeshie :D

ganeshie8 (ganeshie8):

Hahah xD happy new year to you and hardik :)

hartnn (hartnn):

that was the whole point of this question! to wish everyone Happy New Year 2015 ;)

OpenStudy (mathmath333):

oh i see the \(2015\)

OpenStudy (mathmath333):

did'nt notice at first the power was the upcoming year :) ,lol

ganeshie8 (ganeshie8):

I cooked up this problem to be moderately challenging and I'm convinced that I was successful xD Happy new year mathmath :D

OpenStudy (mathmath333):

happy new year to u @ganeshie8 and other friends :)

OpenStudy (freckles):

Sweet question and answers! :) Very flavorful!